# Univariate-Guided Sparse Regression for Biobank-Scale High-Dimensional -omics Data

### v1

@@ Abstract
We present a scalable framework for computing polygenic risk scores (PRS) in high-dimensional genomic settings using the recently introduced Univariate-Guided Sparse Regression (uniLasso). UniLasso is a two-stage penalized regression procedure that leverages univariate coefficients and magnitudes to stabilize feature selection and enhance interpretability. Building on its theoretical and empirical advantages, we adapt uniLasso for application to the UK Biobank, a population-based repository comprising over one million genetic variants measured on hundreds of thousands of individuals from the United Kingdom. We further extend the framework to incorporate external summary statistics to increase predictive accuracy. Our results demonstrate that the adapted uniLasso attains predictive performance comparable to standard Lasso while selecting substantially fewer variants, yielding sparser and more interpretable models. Moreover, it exhibits superior performance in estimating PRS relative to its competitors, such as PRS-CS. Integrating external scores further improves prediction while maintaining sparsity.
@@ 1 Introduction
With rapid advancements in computing power and storage capacity, data collection has become increasingly detailed and large in scale. As a result, scientists routinely encounter high-dimensional data settings, where the number of predictors is much larger than the number of observations ($p\gg n$). Such scenarios are common in fields like genomics and finance, where granular information can be recorded for each individual or transaction. In these regimes, regression has become a central tool for both prediction and variable selection, underpinning a wide range of modern applications. A second defining feature of -omics design matrices in particular, is strong correlation among predictors. In genomics, linkage disequilibrium (LD) creates blocks of SNPs that carry very similar information; in transcriptomics, co-expression and shared regulatory programs induce clusters of correlated genes. A truly causal signal at one locus therefore tends to bring along a cloud of correlated predictors with similar univariate behaviour. Any multivariate method must decide what to do with these groups—either select a sparse set of representatives, or distribute (“smear”) the effect across many correlated variables. Although our main focus is on -omics applications, similar large $n$, larger $p$ structures arise in other domains. In finance, for instance, one may observe long time series or many individual transactions ($n$)) together with hundreds or thousands of asset-, sector-, and macro-level predictors ($p$), many of which move together through shared risk factors or market regimes. High-dimensional sparse regression in such settings faces the same core challenges: correlated predictors, the need to identify a small subset of informative signals, and stringent computational constraints.
Often, a only a much smaller subset of features have a meaningful relationship with the response. Thus a main goal of researchers is to focus on identifying which variables are worth the investment of time and resources to measure and analyze.
@@ 1.1 Review of the lasso
For a design matrix $X\in\mathbb{R}^{n\times p}$ and response vector $y\in\mathbb{R}^{n}$, Lasso solves the following minimization problem:
@@
$\lambda$ is a regularization parameter that determines the number of features identified in the solution set. This optimization is typically solved over a grid of $\lambda$-values, with the optimal one chosen via cross-validation or by performance of the algorithm on a hold-out validation set. Equivalently, we can formulate the Lasso problem in the following way:
@@ 1.2 Optimization Approaches for the Lasso
A variety of computational tools have been developed to efficiently solve Lasso problems. glmnet [Friedman et al., 2010] is among the most widely cited R packages, capable of computing entire solution paths for a range of generalized linear models. snpnet [Qian et al., 2020] is another R package designed specifically for large-scale, high-dimensional Lasso regression in genomics applications. snpnet stands out as one of the most competitive solvers, achieving high predictive accuracy in such settings.
@@
Adelie [Yang and Hastie, 2024] is a novel Python package designed for efficiently solving lasso and group lasso problems. This package can handle large-scale inputs with greater computational efficiency than traditional solvers, and the package provides a suite of generalized linear model (GLM) classes. Benchmarks against other state-of-the-art solvers, such as glmnet, demonstrate that Adelie achieves faster speeds than other group lasso packages with equivalent accuracy. Additionally, Adelie offers flexible parameter tuning, warm-start capabilities, and easy integration with standard Python data science workflows, making it powerful for high-dimensional regression tasks.
Another notable feature of Adelie is its ability to efficiently perform regression analyses on genome-wide association study (GWAS) datasets. To handle the scale and complexity of such data, Adelie introduces specialized file formats optimized for both memory efficiency and the speed of matrix operations. This design allows the package to process large-scale genotype and phenotype matrices that would otherwise exceed standard memory limits, enabling scalable analyses of GWAS data.
@@ 1.3 The UK Biobank Database
Regression provides a natural framework for studying the UK Biobank data, allowing researchers to quantify relationships between SNPs and phenotypes. Packages such as Adelie enable analyses to be performed efficiently on such large-scale datasets. By incorporating regularization techniques, i.e. Lasso, regression can produce sparse solutions, effectively reducing the set of over a million SNPs to a smaller subset most likely to be associated with the response. This not only aids in interpretability, but also improves predictive performance by mitigating overfitting in high-dimensional settings.
@@ 1.4 Contributions
In this paper, we propose a novel method for analyzing ultra-high-dimensional -omics data that directly fits a sparse multivariate model on individual-level data, rather than relying on the standard two-step GWAS-plus-PRS pipeline. Our work focuses on constructing polygenic risk scores (PRS) from biobank-scale SNP data, where both sparsity and interpretability of selected variants are crucial. We present an adaptation of Univariate-Guided Sparse Regression (uniLasso) [Chatterjee et al., 2025] tailored for this challenging data regime. Empirically, uniLasso achieves predictive performance comparable to existing methods but with  40% fewer predictors, significantly improving the interpretability of the results. Additionally, we consider the integration of external data sources to further enhance predictive accuracy. In biomedical research, privacy regulations often prevent the sharing of raw data, so researchers sometimes may only provide summary statistics rather than entire data sets.
Following the approach proposed in Chatterjee et al. [2025], we demonstrate how to incorporate external scores from independent datasets into the uniLasso pipeline, bolstering predictions. We show how these external scores can be combined with individual-level data in a single sparse multivariate model, rather than as a separate post-GWAS step.
@@ 2.1 Overview of the uniLasso Algorithm
The uniLasso procedure, as outlined in Chatterjee et al. [2025], assumes the same supervised learning framework as Lasso. For a design matrix $X\in\mathbb{R}^{n\times p}$ and response vector $y\in\mathbb{R}^{n}$, the uniLasso procedure is defined as follows:
@@
    For $j=1,2,\ldots,p$ compute the univariate intercepts and slopes $(\hat{\beta}_{0j},\hat{\beta}_{j})$ and their leave-one-out (LOO) counterparts $(\hat{\beta}_{0j}^{-i},\hat{\beta}_{j}^{-i}),\;i=1,\ldots,n$. Compute the LOO fitted values for all $i$ and $j$ and aggregate them in a new $n\times p$ feature matrix $F.$
@@
    Fit a non-negative Lasso with an intercept and no standardization to target $y$ and the univariate LOO predictions as features:
    $\operatorname*{argmin}_{\theta}\quad\left\{\frac{1}{n}\sum_{i=1}^{n}\Bigl(y_{i}-\theta_{0}-\sum_{j=1}^{p}(\hat{\beta}_{0j}^{-i}+\hat{\beta}_{j}^{-i}x_{ij})\,\theta_{j}\Bigr)^{2}+\lambda\sum_{j=1}^{p}\theta_{j}\right\}$
    $\text{subject to}\quad\theta_{j}\geq 0,\quad\forall j>0.\text{ Select $\lambda$ through cross validation.}$
@@
The steps taken in this procedure serve several purposes. First, it ensures that the signs of the final coefficients agree with the signs of the univariate coefficients (or they are zero). This increases the interpretability of the result, as it excludes features whose coefficient sign changes between the univariate and multivariate model. In addition, features with larger univariate coefficients will also tend to have larger coefficients in the final model. Moreover, using LOO fitted values rather than raw univariate estimates or the original features amounts to feeding cross-validated univariate predictions into the Lasso in Step 2. This forces the selection step to rely on out-of-sample predictive signal rather than in-sample noise, which empirically leads to sparser models and improved generalization. [Chatterjee et al., 2025].
As demonstrated in the uniLasso paper, this method has exhibited excellent performance, producing sparser solutions than classical techniques. However, many challenges emerge when applying the algorithm in the high-dimensional regime. In its original formulation, however, uniLasso is not immediately scalable to biobank-sized datasets, which prevents a direct comparison with genomics tools such as snpnet.
As discussed above, one challenge of ultra-high-dimensional data analysis is the need to store the immense feature matrix. In computation, Step 1 of the uniLasso pipeline actually requires the user to remember another $(n\times p)$ matrix of the LOO fitted values. While SNP matrices are typically sparse (many of the entries are 0), the LOO matrix $F$ is not. Thus the computer must be capable of storing another feature matrix which could take up over 1500 GB of memory.
@@ 2.2 Details of the Omics-adapted uniLasso Algorithm
We utilize Sherlock, a high-performance computing (HPC) cluster at Stanford University, to engineer the uniLasso pipeline using the Adelie Python package. Our compute instances use a cluster of Intel(R) Xeon(R) E5-2640 v4 CPUs with 64 total cores. This setup enables efficient allocation of memory and computational resources necessary to sustain the full analysis. When calculating the leave-one-out (LOO) fits and univariate coefficients, it is important to note that each computation depends only on individual columns of the design matrix.
For a given SNP column $x_{i}$, let $x_{c}=x_{i}-\bar{x}$, where $\bar{x}$ denotes the column mean. Similarly, define $y_{i}$ as the response and $y_{c}=y_{i}-\bar{y}$ as its centered counterpart. The univariate slope $u_{1}$ and intercept $u_{0}$ are then given by:
$u_{1}=\frac{x_{c}^{T}y_{c}}{x_{c}^{T}x_{c}},\hskip 56.9055ptu_{0}=\bar{y}-u_{1}\cdot\bar{x}.$
The LOO fitted values $L$ for the same column $x_{i}$ can be computed as follows, where $H_{ii}$ denotes the $i^{th}$ diagonal element of the hat matrix $H$:
$L=y_{i}-\frac{y_{i}-(u_{0}+u_{1}x_{i})}{1-H_{ii}},\quad\text{where }H_{ii}=\frac{1}{n}+\frac{x_{c}\odot x_{c}}{x_{c}^{T}x_{c}}.$
Using vectorized formulas we parallelized the computations, which occur in an embarrassingly parallel manner. This allowed us to fully exploit CPU resources and parallel computing to markedly accelerate computation.
In the standard uniLasso algorithm, the regularization parameter in Step 2 is typically chosen via $k$-fold cross validation. However, once again given the scale of our data, this approach is computationally infeasible. Instead, we employed a train-validation-test split (70/10/20) to efficiently tune the regularization strength. Specifically, we first fit the model on the training set over a grid of $\lambda$ values. Using the validation set, we then identified the value of $\lambda$ that achieved the best predictive performance. Next, we refit the model on the combined training and validation sets with the chosen parameter. Finally, we evaluated the resulting model on the held-out test set to obtain the optimal coefficients.
@@
    $S_{xx}=x_{c}^{T}x_{c},\quad\text{where }x_{c}=x-\bar{x}.$
@@ 2.3 Omics-adapted uniLasso Algorithm for Binary Response Variables
For uniLasso, recall that the algorithm requires LOO fitted values and univariate coefficients. Unlike the case with Gaussian loss, however, there are no closed-form solutions for these expressions. However, in practice, logistic regression is typically fit via iteratively reweighted least squares (IRLS). Given the fitted linear predictor vector $\eta_{j}^{(\ell)}$ at iteration $\ell$, one forms a working response vector $z_{j}^{(\ell)}$ that depends on $\eta_{j}^{(\ell)}$, $y$, and the GLM family, along with an observation weight vector $w_{j}^{(\ell)}$. The updated model $\eta_{j}^{(\ell+1)}$ is then obtained by performing weighted least squares of $z_{j}^{(\ell)}$ on $x_{j}$ with weights $w_{j}^{(\ell)}$. In our implementation, we perform two IRLS iterations per feature.
@@
Finally, we apply the same regularization and MAF filtering when running the binary uniLasso algorithm. Aside from the modifications required for the logistic loss and the corresponding univariate values, the majority of the uniLasso pipeline remains nearly identical to the Gaussian case.
@@ 3 UK Biobank Analysis
The UK Biobank [Bycroft et al., 2018] is a large prospective cohort study containing extensive genetic and phenotypic data from approximately 500,000 middle-aged individuals across the United Kingdom. In this analysis, we focus on modeling the relationship between individual genotype sequences and their corresponding phenotypic traits. Using uniLasso, we aim to identify SNPs that are most strongly associated with the phenotype of interest. Our proposed adaptation enables model fitting at full scale and yields a reduced active variable set compared to alternative methods. This smaller collection of selected SNPs enhances interpretability while maintaining comparable levels of variance explained.
We restricted our analysis to a subset of 336,442 White British individuals from the full UK Biobank dataset who satisfy the population stratification criteria described in DeBoever et al. [2018]. The data were randomly partitioned into training ($70\%$), validation ($10\%$), and test ($20\%$) subsets. Each individual has 1,080,968 measured genetic variants, where each variant is encoded by one of three levels: 0 for homozygous major alleles, 1 for heterozygous alleles, and 2 for homozygous minor alleles. Missing genotypes were imputed using the mean of the observed values. In addition to genotype data, covariates such as age, sex, and 10 precomputed principal components of the SNP matrix were included to account for population structure. Among the thousands of measured phenotypes in this dataset, the phenotypes we used to evaluate our method were standing height, body mass index (BMI), coronary heart disease (CHD), and asthma. CHD and asthma are binary responses, where 0 indicates a lack of the disease and 1 delineates individuals who do suffer from the disease.
@@
Table 1: Number of non-zero coefficients for each phenotype measured across all Lasso-based algorithms. Compared to Lasso, uniLasso selects 38% fewer SNPs for height and 43% fewer for BMI, while only reducing $R^{2}$ by 0.006 and 0.016, respectively.
@@
where $\bar{y}_{0}$ is the average response of the training data. In this setting, the PRS for individual $i$ is the linear predictor $\widehat{\mathrm{PRS}}_{i}=x_{i}^{\top}\hat{\beta}$, where $x_{i}$ denotes that individual’s vector of SNP genotypes. This metric represents the proportion of variance explained by the covariates. Ranging from 0 to 1, higher values represent better explanatory power by the model. We evaluate this criteria for all models fit on continuous response variables.
@@
For binary responses, we assess predictive performance using the receiver operating characteristic (ROC) curve, which illustrates the trade-off between true and false positive rates across different classification thresholds. The area under the ROC curve (AUC) reflects how well the model distinguishes between positive and negative outcomes, with higher values corresponding to improved performance. For phenotypes with binary outcomes, we evaluate the AUC for all regression methods used.
@@
We perform Lasso and uniLasso on all of the aforementioned phenotypes to have a sense of any potential advantages. For the continuous phenotypes (height and BMI) there is a consistent trend: Lasso outperforms uniLasso by a small margin but uses a significantly larger number of coefficients in its solution. The small loss in $R^{2}$ is worthwhile for uniLasso when considering the nearly $40\%$ reduction in active set size. Figure 1 illustrates the difference in their performance on held out test sets while Table 1 displays the number of non-zero coefficients at the optimal $\lambda$ value for each method tested on each phenotype.
@@
Figure 1: Comparison of the test set predictive performance of the different polygenic risk score (PRS) methods with refitting on the training and the validation set. All regressions were performed with Adelie in Python. $R^{2}$ are evaluated for the continuous phenotypes (height and BMI), and AUC evaluated for the binary phenotypes (asthma and CHD).
@@
In the binary case, uniLasso and Lasso perform much more similarly when measured in AUC, with uniLasso even outperforming Lasso on Asthma. Otherwise, the difference in test error between Lasso and uniLasso does not exceed 0.016, an encouragingly small margin. Excitingly, as also seen with the continuous responses, uniLasso maintains a significantly sparser solution set than Lasso. For CHD, uniLasso is over 50% sparser than Lasso and has lower test error by only 0.001. UniLasso is consistently the best performing under this sparsity metric. This reduced solution set in conjunction with the alignment of signs between univariate and multivariate coefficients reinforces the merits of uniLasso. Section 4 goes into more detail about an alternative formulation of uniLasso that potentially achieves a more optimal tradeoff.
@@
Table 2 summarizes the number of sign changes observed between the univariate coefficients and the non-zero coefficients in the final Lasso model. This metric reflects the interpretability of the multivariate (Lasso) model: solutions with sign changes indicate that the multivariate fit has flipped the direction of association for certain variables. Such reversals are difficult to interpret because they obscure the role each variable plays in predicting disease susceptibility (or any other phenotype); the meaning of a coefficient’s sign becomes context-dependent, which is particularly ambiguous when dealing with SNPs. As shown in the table, the binary models exhibit far fewer sign changes than the continuous ones, which is expected given their smaller active sets. A built-in advantage of uniLasso is the lack of sign changes regardless of the type of response variable.
Number of Sign Changes in Lasso
Phenotype Num. Sign Changes
Height 6,698
BMI 103
CHD 2
Asthma 2
@@
Table 2: Number of coefficient sign changes between the univariate and multivariate models for each phenotype. UniLasso, by construction, eliminates all such sign reversals; every non-zero coefficient retains the sign of its univariate effect.

### v4

# Univariate-Guided Sparse Regression for Biobank-Scale High-Dimensional Omics Data

@@ Abstract
We present a scalable framework for computing polygenic risk scores (PRS) in high-dimensional genomic settings using the recently introduced Univariate-Guided Sparse Regression (uniLasso). UniLasso is a two-stage penalized regression procedure that leverages univariate coefficients and magnitudes to stabilize feature selection and enhance interpretability. Building on its theoretical and empirical advantages, we adapt uniLasso for application to the UK Biobank, a population-based repository comprising over one million genetic variants measured on hundreds of thousands of individuals from the United Kingdom. We further extend the framework to incorporate external summary statistics to increase predictive accuracy. Our results demonstrate that uniLasso attains predictive performance comparable to standard Lasso while selecting substantially fewer variants, yielding sparser and more interpretable models. Moreover, it exhibits superior performance in estimating PRS relative to its competitors, such as PRS-CS. Integrating external scores further improves prediction while maintaining sparsity.
@@ 1 Introduction
With rapid advancements in computing power and storage capacity, data collection has become increasingly detailed and large in scale. As a result, scientists routinely encounter high-dimensional data settings, where the number of predictors is much larger than the number of observations ($p\gg n$). Such scenarios are common in fields like genomics and finance, where granular information can be recorded for each individual or transaction. In these regimes, regression has become a central tool for both prediction and variable selection, underpinning a wide range of modern applications. A second defining feature of -omics design matrices, in particular, is strong correlation among predictors. In genomics, linkage disequilibrium (LD) creates blocks of SNPs that carry very similar information; in transcriptomics, co-expression and shared regulatory programs induce clusters of correlated genes. A truly causal signal at one locus therefore tends to bring along a cloud of correlated predictors with similar univariate behaviour. Any multivariate method must decide what to do with these groups—either select a sparse set of representatives, or distribute (“smear”) the effect across many correlated variables. Although our main focus is on -omics applications, similar large $n$, larger $p$ structures arise in other domains. In finance, for instance, one may observe long time series or many individual transactions ($n$) together with hundreds or thousands of asset-, sector-, and macro-level predictors ($p$), many of which move together through shared risk factors or market regimes. High-dimensional sparse regression in such settings faces the same core challenges: correlated predictors, the need to identify a small subset of informative signals, and stringent computational constraints.
Often, only a much smaller subset of features have a meaningful relationship with the response. Thus a main goal of researchers is to focus on identifying which variables are worth the investment of time and resources to measure and analyze.
@@ 1.1 Review of the Lasso
For a model matrix $X\in\mathbb{R}^{n\times p}$ and response vector $y\in\mathbb{R}^{n}$, Lasso solves the following minimization problem:
@@
$\lambda$ is a regularization parameter that determines the number of features identified in the solution set. In practice, we evaluate the objective over a grid of $\lambda$ values and then select the best $\lambda$ using either cross-validation or performance on a held-out validation set. Equivalently, we can formulate the Lasso problem in the following way:
@@ 1.2 Optimization Approaches for the Lasso
A variety of computational tools have been developed to efficiently solve Lasso problems. glmnet [Friedman et al., 2010] is among the most widely cited R packages, capable of computing entire solution paths for a range of generalized linear models. snpnet [Qian et al., 2020] is another R package designed specifically for large-scale, high-dimensional Lasso regression in genomics applications. snpnet stands out as one of the most competitive solvers, combining strong-rule screening with efficient KKT batching to scale to millions of SNPs. This approach yields high predictive accuracy in large-scale genomic analyses.
@@
Adelie [Yang and Hastie, 2024] is a novel Python package (with a version in R as well) designed for efficiently solving lasso and group lasso problems. This package can handle large-scale inputs with greater computational efficiency than traditional solvers, and the package provides a suite of generalized linear model (GLM) classes. Benchmarks against other state-of-the-art solvers demonstrate that Adelie achieves faster speeds than other group lasso packages with equivalent accuracy. On Lasso and EN problems, it performs comparably to the glmnet R package. Additionally, Adelie offers flexible parameter tuning, warm-start capabilities, and easy integration with standard Python data science workflows, making it powerful for high-dimensional regression tasks.
Another notable feature of Adelie, similar to snpnet, is its ability to efficiently fit lasso and EN models on genome-wide association study (GWAS) datasets in Python. To handle the scale and complexity of such data, Adelie introduces specialized file formats optimized for both memory efficiency and the speed of matrix operations. This design allows the package to process large-scale genotype and phenotype matrices that would otherwise exceed standard memory limits, enabling scalable analyses of GWAS data.
@@ 1.3 The UK Biobank Database
Regression provides a natural framework for studying the UK Biobank data, allowing researchers to quantify relationships between SNPs and phenotypes. Packages such as Adelie and snpnet enable analyses to be performed efficiently on such large-scale datasets. By incorporating regularization techniques, i.e. Lasso, regression can produce sparse solutions, effectively reducing the set of over a million SNPs to a smaller subset most likely to be associated with the response. This not only aids in interpretability, but also improves predictive performance by mitigating overfitting in high-dimensional settings.
@@ 1.4 Contributions
In this paper, we implement a novel method for analyzing ultra-high-dimensional -omics data that directly fits a multivariate model on individual-level data, rather than relying on the standard two-step GWAS-plus-PRS pipeline. We present an adaptation of Univariate-Guided Sparse Regression (uniLasso) [Chatterjee et al., 2025] tailored for biobank-scale SNP data. Empirically, uniLasso achieves predictive performance comparable to existing methods but with  40% fewer predictors, significantly improving the interpretability of the results.
Additionally, we consider the integration of external data sources to further enhance predictive accuracy. In biomedical research, privacy regulations often prevent the sharing of raw data, so researchers sometimes may only provide summary statistics rather than entire data sets. Following the approach proposed in Chatterjee et al. [2025], we demonstrate how to incorporate external scores from independent datasets into the uniLasso pipeline, bolstering polygenic risk score (PRS) predictions. We show how these external scores can be combined with individual-level data in a single sparse multivariate model, rather than as a separate post-GWAS step.
@@ 2.1 Overview of the uniLasso Algorithm
The uniLasso procedure, as outlined in Chatterjee et al. [2025], assumes the same supervised learning framework as Lasso. For a model matrix $X\in\mathbb{R}^{n\times p}$ and response vector $y\in\mathbb{R}^{n}$, the uniLasso procedure is defined as follows:
@@
    For $j=1,2,\ldots,p$ compute the univariate intercepts and slopes $(\hat{\beta}_{0j},\hat{\beta}_{j})$ and the leave-one-out (LOO) fitted values
    $\hat{y}^{-i}_{ij}\;=\;\hat{\beta}_{0j}^{-i}+\hat{\beta}_{j}^{-i}x_{ij},\qquad i=1,\ldots,n,\;j=1,\ldots,p,$
    and collect them in a new $n\times p$ feature matrix $F$ with entries $F_{ij}=\hat{y}^{-i}_{ij}$.
@@
    Fit a non-negative lasso with an intercept and no standardization to target $y$ using the LOO univariate predictions in $F$ as features:
    $\operatorname*{argmin}_{\theta}\quad\left\{\frac{1}{n}\sum_{i=1}^{n}\Bigl(y_{i}-\theta_{0}-\sum_{j=1}^{p}F_{ij}\,\theta_{j}\Bigr)^{2}+\lambda\sum_{j=1}^{p}\theta_{j}\right\}$
    $\text{subject to}\quad\theta_{j}\geq 0,\quad j=1,\ldots,p.\ \text{Select $\lambda$ through cross validation.}$
@@
The steps taken in this procedure serve several purposes. First, it ensures that the signs of the final coefficients agree with the signs of the univariate coefficients (or they are zero). This increases the interpretability of the result, as it excludes features whose coefficient sign changes between the univariate and multivariate model. In addition, features with larger univariate coefficients will also tend to have larger coefficients in the final model, since they are given an advantage in the penalization of $\theta$. Moreover, using LOO fitted values rather than raw univariate estimates or the original features amounts to feeding cross-validated univariate predictions into the Lasso in Step 2. This forces the selection step to rely on out-of-sample predictive signal rather than in-sample noise, which empirically leads to sparser models and improved generalization [Chatterjee et al., 2025]. It also justifies the use of cross validation.
As demonstrated in the uniLasso paper, this method has exhibited excellent performance, producing sparser solutions than classical techniques. In its original formulation, however, uniLasso is not immediately scalable to biobank-sized datasets, which prevents a direct comparison with genomics tools such as snpnet.
One challenge posed by uniLasso in ultra-high-dimensional data analysis is the need to store the immense feature matrix. In computation, Step 1 of the uniLasso pipeline actually requires the user to remember another $(n\times p)$ matrix of the LOO fitted values. While SNP matrices are typically sparse (many of the entries are 0), the LOO matrix $F$ is not. Thus the computer must be capable of storing another feature matrix which could take up over 1500 GB of memory.
@@ 2.2 Details of the Omics-adapted uniLasso Algorithm
We use Sherlock, a high-performance computing (HPC) cluster at Stanford University, to engineer the uniLasso pipeline using the Adelie Python package. Our compute instances use a cluster of Intel(R) Xeon(R) E5-2640 v4 CPUs with 64 total cores. This setup enables efficient allocation of memory and computational resources necessary to sustain the full analysis. When calculating the leave-one-out (LOO) fits and univariate coefficients, it is important to note that each computation depends only on individual columns of the design matrix.
For a given SNP column $x$, let $\tilde{x}=x-\bar{x}$, where $\bar{x}$ denotes the column mean. Similarly, define $y$ as the response and $\tilde{y}=y-\bar{y}$ as its centered counterpart. The univariate slope $u_{1}$ and intercept $u_{0}$ are then given by:
$u_{1}=\frac{\tilde{x}^{\top}\tilde{y}}{\|\tilde{x}\|^{2}},\hskip 56.9055ptu_{0}=\bar{y}-u_{1}\cdot\bar{x}.$
The vector of LOO fitted values $\vec{f}$ corresponding to the column $x$ are defined as follows, where $H_{ii}$ denotes the $i^{th}$ diagonal element of the hat matrix $H$:
$y_{i}-f_{i}=\frac{y_{i}-(u_{0}+u_{1}x_{i})}{1-H_{ii}},\quad\text{where }H_{ii}=\frac{1}{n}+\frac{\tilde{x_{i}}^{2}}{\|\tilde{x}\|^{2}}.$
Using vectorized formulas suggested in Chatterjee et al. [2025] we parallelized the computations, which occur in an embarrassingly parallel manner. This allowed us to fully exploit CPU resources and parallel computing to markedly accelerate computation.
In the standard uniLasso algorithm, the regularization parameter in Step 2 is typically chosen via $k$-fold cross validation. However, once again given the scale of our data, this is computationally infeasible. Instead, following the approach of Qian et al. [2020], we employed a train-validation-test split (70/10/20) to efficiently tune the regularization strength. Specifically, we first fit the model on the training set over a grid of $\lambda$ values. Using the validation set, we then identified the value of $\lambda$ that achieved the best predictive performance. Next, we refit the model on the combined training and validation sets with the chosen parameter. Finally, we evaluated the resulting model on the held-out test set to obtain the optimal coefficients.
@@
    $S_{xx}=\tilde{x}^{\top}\tilde{x},\quad\text{where }\tilde{x}=x-\bar{x}.$
@@ 2.3 Omics-adapted uniLasso Algorithm for Binary Response Variables
For uniLasso, recall that the algorithm requires LOO fitted values and univariate coefficients. Unlike the case with Gaussian loss, there are no closed-form solutions for these expressions. However, in practice, logistic regression is typically fit via iteratively reweighted least squares (IRLS). We follow the procedure described in Chatterjee et al. [2025], summarized below:
Given the fitted linear predictor vector $\eta_{j}^{(\ell)}$ at iteration $\ell$, one forms a working response vector $z_{j}^{(\ell)}$ that depends on $\eta_{j}^{(\ell)}$, $y$, and the GLM family, along with an observation weight vector $w_{j}^{(\ell)}$. The updated model $\eta_{j}^{(\ell+1)}$ is then obtained by performing weighted least squares of $z_{j}^{(\ell)}$ on $x_{j}$ with weights $w_{j}^{(\ell)}$. In our implementation, we perform two IRLS iterations per feature.
@@
Uniquely, we apply the same regularization and MAF filtering when running the binary uniLasso algorithm. Aside from the modifications required for the logistic loss and the corresponding univariate values, the majority of the uniLasso pipeline remains nearly identical to the Gaussian case.
@@ 3 UK Biobank Analysis
The UK Biobank [Bycroft et al., 2018] is a large prospective cohort study containing extensive genetic and phenotypic data from approximately 500,000 middle-aged individuals across the United Kingdom. In this analysis, we focus on modeling the relationship between individual genotype sequences and their corresponding phenotypic traits. Using uniLasso, we aim to identify SNPs that are most informative for predicting the phenotype of interest. Our proposed adaptation enables model fitting at full scale and yields a reduced active variable set compared to alternative methods. This smaller collection of selected SNPs enhances interpretability while maintaining comparable levels of variance explained.
We restricted our analysis to a subset of 336,442 White British individuals from the full UK Biobank dataset who satisfy the population stratification criteria described in DeBoever et al. [2018]. The data were randomly partitioned into training ($70\%$), validation ($10\%$), and test ($20\%$) subsets. Each individual has 1,080,968 measured genetic variants, where each variant is encoded by one of three levels: 0 for homozygous major alleles, 1 for heterozygous alleles, and 2 for homozygous minor alleles. Missing genotypes were imputed using the mean of the observed values. In addition to genotype data, covariates such as age, sex, and 10 precomputed principal components of the SNP matrix were included to account for population structure. Among the thousands of measured phenotypes in this dataset, the phenotypes we used to evaluate our method were standing height, body mass index (BMI), coronary heart disease (CHD), and asthma. CHD and asthma are binary responses, where 1 indicates disease and 0 no disease.
@@
Table 1: Number of non-zero coefficients for each phenotype measured across all Lasso-based algorithms. Compared to Lasso, uniLasso selects 38% fewer SNPs for height and 43% fewer for BMI, while only reducing $R^{2}$ by 0.006 and 0.016, respectively. uniLasso with external scores (uniLasso ES) is introduces in Section 4.
@@
where $\bar{y}_{0}$ is the average response of the training data and 1 is the $n$-vector of all ones. The PRS for individual $i$ is the linear predictor $\widehat{\mathrm{PRS}}_{i}=x_{i}^{\top}\hat{\beta}$, where $x_{i}$ denotes that individual’s vector of SNP genotypes. In this formulation, covariates enter only through the evaluation model used to predict the phenotype itself. The resulting $R^{2}$ quantifies the proportion of variance explained by this combined model. Ranging from 0 to 1, higher values represent better explanatory power by the model.
@@
For binary responses, we assess predictive performance using the receiver operating characteristic (ROC) curve, which illustrates the trade-off between true and false positive rates across different classification thresholds. The area under the ROC curve (AUC) reflects how well the model distinguishes between positive and negative outcomes, with values ranging from 0.5 (no discriminatory ability) to 1 (perfect discrimination). For phenotypes with binary outcomes, we evaluate the AUC for all regression methods used.
@@
We perform Lasso and uniLasso on all of the aforementioned phenotypes to have a sense of any potential advantages. For the continuous phenotypes (height and BMI) there is a consistent trend: Lasso outperforms uniLasso by a small margin but uses a significantly larger number of coefficients in its solution. The small loss in $R^{2}$ is worthwhile for uniLasso when considering the nearly $40\%$ reduction in active-set size. Figure 1 illustrates the difference in their performance on held out test sets while Table 1 displays the number of non-zero coefficients at the optimal $\lambda$ value for each method tested on each phenotype.
@@
Figure 1: Comparison of the test set predictive performance of the different polygenic risk score (PRS) methods with refitting on the training and the validation set. All regressions were performed with Adelie in Python. Test-set $R^{2}$ is the metric evaluated for the continuous phenotypes (height and BMI), and AUC is evaluated for the binary phenotypes (asthma and CHD).
@@
In the binary case, uniLasso and Lasso perform much more similarly as measured by AUC, with uniLasso slightly outperforming Lasso on Asthma. Otherwise, the difference in test error between Lasso and uniLasso does not exceed 0.016, an encouragingly small margin. Excitingly, as also seen with the continuous responses, uniLasso maintains a significantly sparser solution set than Lasso (See Table 1). For CHD, uniLasso is over 50% sparser than Lasso and has lower test error by only 0.001. This reduced solution set in conjunction with the alignment of signs between univariate and multivariate coefficients reinforces the merits of uniLasso. Section 4 goes into more detail about an alternative formulation of uniLasso that potentially achieves a more optimal tradeoff.
@@
Table 2 summarizes the number of sign changes observed between the univariate coefficients and the non-zero coefficients in final Lasso model. This metric is a reflection of the interpretability of the multivariate (Lasso) model: sign changes indicate that the multivariate fit has flipped the direction of association for certain variables. Such reversals are difficult to interpret because the coefficient’s sign becomes dependent on the presence of other variables. As shown in the table, the binary models exhibit far fewer sign changes than the continuous ones, which is expected given their smaller active sets. A built-in advantage of uniLasso is the lack of sign changes regardless of the type of response variable.
Number and Percentage of Sign Changes in Lasso
Phenotype # Sign Changes % Sign Changes
Height 6,698 12%
BMI 103 0.3%
CHD 2 $\leq$ 0.01%
Asthma 2 $\leq$ 0.01%
@@
Table 2: Number of coefficient sign changes between the univariate and multivariate models for each phenotype and the percentage of coefficients that switched. UniLasso, by construction, eliminates all such sign reversals; every non-zero coefficient retains the sign of its univariate effect.